OpenStudy (anonymous):

Please help me understand....? Which of the following represents the area of a rectangle whose length is x - 7 and whose width isx + 12? Answer choices: x2 + 5x - 84 x2 - 84 x2 - 5x - 84 x2 - 84x + 5s

OpenStudy (anonymous):

the area of a rectangle is w*l, and you are given w = x+12, l=x-7, the equation for the area will be w*l = (x+12)(x-7), then you expand it

OpenStudy (baldymcgee6):